1. Assume that you have solid NaCl and all the glassware in your lab drawer available to you. What mass of sodium chloride would you need to prepare 10.0 mL of a 1.0 M NaCl(aq) solution.

2. Using the solution you prepared in #1, how many mL would you need to prepare a 2.5 mL 0.10 M NaCl solution?
Answer #1

NACl

M = 1

V = 10 ml or 0.01 L

n = M*V = 1*10/1000 = 0.01 mol o fNaCl

MW = 58

mass = mol*MW = 0.01*58 = 0.58 grams of NaCl

2)

ml for

V = 2.5 ml of a 0.10

M1V1 = M2V2

V1 = M2/M1 * V2

V1 = 0.1/1 * 2.5 = 0.25 ml are needed of original mix